OpenStudy (anonymous):

What is the surface area of the box if it is scaled up by a factor of 10? Length: 9in. Width: 5in. Height: 3in.

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

will fan and medal just please help

OpenStudy (anonymous):

If it is scaled up by 10, then all the side lengths are multiplied by 10.

OpenStudy (anonymous):

Do you know how to find the surface area?

OpenStudy (anonymous):

yeahh and its 174

OpenStudy (anonymous):

\[ S = 2lw+2lh+2wh \]

OpenStudy (anonymous):

If you scale it up by 10 \[ S' = 2l'w'+2l'h'+2w'h' = 2(10l)(10w)+2(10l)(10h)+2(10w)(10h)\\ = 100(2lw + 2lh+2wh) = 100S \]

OpenStudy (anonymous):

So the surface area is increased by a factor of \(10^2=100\)

OpenStudy (anonymous):

ok so I just multiply each by 10??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

nvm, I got the answer but I need help with one more

OpenStudy (anonymous):

What is the volume of the box if it is scaled down by a factor of 1/10?

OpenStudy (anonymous):

The volume would be scaled down by \(1/10^3 = 1/1000\).

OpenStudy (anonymous):

ok so what do I do??

OpenStudy (anonymous):

please help me